# On determinant of checkerboard colorable virtual knots

### v1

@@ 3. Ascending and descending polynomials
We review basic definitions according to [7]. An arrow diagram is a based chord diagram with oriented chords (called arrows). A homomorphism $\varphi$ from $A$ to $G$ is an orientation-preserving homeomorphism of the circle of $A$ to the circle of $G$ which maps the basepoint to the basepoint and induces an injective map of arrows of $A$ to arrows of $G$ respecting the orientation of the arrows.
@@
For a non-negative integer $n$, the Conway combination $\mathfrak{C}_{2n}$ (resp. $\mathfrak{C}^{\prime}_{2n}$) is defined to be the sum of one-component ascending (resp. descending) arrow diagrams with $2n$ arrows on a circle (see [7, Definition 4.3] for the definitions of one-component descending and ascending arrow diagrams). Similarly, $\mathfrak{C}_{2n+1}$ (resp. $\mathfrak{C}^{\prime}_{2n+1}$) is defined to be the sum of one-component ascending (resp. descending) diagrams with $12n+$ 1 arrows on two circles.
@@
The polynomials $\nabla_{\mathrm{asc}}(G)(z)$ and $\nabla_{\mathrm{des}}(G)(z)$ depend on the choice of a basepoint of $G$ in general. For instance, one can observe it for the Gauss diagram in Figure 1. This is because the virtual knot in Figure 1 is not (mod 00) almost classical, which implies that the assumption of Theorems 1.1 and 4.7 is essential in their proofs.
@@
Let $K$ be the (almost classical) virtual knot $6.87548$ in Green’s table [12]. For the Gauss diagram in [9, Figure 14] (with an arbitrary basepoint), the coefficients of $z^{2}$ in $\nabla_{\mathrm{asc}}(G)(z)$ and in $\nabla_{\mathrm{des}}(G)(z)$ are $-2$. On the other hand, concerning determinant and Arf invariant in [3], one can see $\det(K,F)=5$ and $\operatorname{Arf}(q_{K,F})=1$. Indeed, these are computed from a Seifert matrix
@@
written in [9, Example (6.87548)]. See also [3, Remark 3.5].
In [7, Lemma 5.1 and (5.3)], skein relations among pairings are shown for classical links. In their proofs, any property of classical knots is not used, and hence we obtain the following lemmas as extensions to virtual links.
@@
Let $D_{+}$, $D_{-}$, $D_{0}$ be a skein triple drawn in Figure 4 such that $D_{\pm}$ have one component and $D_{0}$ has two components. Let $G_{+}$, $G_{-}$, $G_{0}$ be the corresponding Gauss diagrams. Then, for $n\geq 1$,
@@
Let $D_{+}$, $D_{-}$, $D_{0}$ be a skein triple drawn in Figure 4 such that $D_{\pm}$ have two components and $D_{0}$ has one component. Let $G_{+}$, $G_{-}$, $G_{0}$ be the corresponding Gauss diagrams. Then, for $n\geq 0$,
@@
The next notion is an extension of [17] to virtual knots, which will be used in the proof of Theorem 4.5.
@@
Let $D$ be a diagram of an oriented virtual knot with a basepoint. The warping degree $d(D)$ of $D$ is defined to be the number of crossings at which we first go through the under-arcs when we travel from the basepoint along the orientation of $D$. Moreover, the warping degree $d(G)$ of a based Gauss diagram $G$ is defined to be that of a virtual link diagram whose Gauss diagram is $G$.
A diagram $D$ is said to be descending if $d(D)=0$. It is shown in [16, Proposition 2.2] that a descending diagram is welded equivalent to the trivial one. This fact will be used in the proof of Lemma 4.3.
@@ 4. Main results
First, $\langle\mathfrak{C}_{2n},G_{b}\rangle=0$ is obvious by definition. Next, since $d(G)=0$, every descending subdiagram of $G$ is of the form in Figure 7. This is not one-component, and thus $\langle\mathfrak{C}^{\prime}_{2n},G_{b}\rangle=0$. ∎
@@
Let $G_{b}$ be a based Gauss diagram of mod $p$ almost classical oriented virtual knot and let $\mathsf{a}$ be an arrow. The $\mathsf{a}$ divides the circle into two arcs and suppose that the tail of any arrow intersecting with $\mathsf{a}$ is attached to the arc containing $b$. Let $H_{b}$ be the $2$-component based Gauss diagram obtained by smoothing $G_{b}$ along $\mathsf{a}$. Then, $\langle\mathfrak{C}_{1},H_{b}\rangle\equiv\langle\mathfrak{C}^{\prime}_{1},H_{b}\rangle\equiv 0\bmod p$ and $\langle\mathfrak{C}_{2n-1},H_{b}\rangle=0$ for $n\geq 2$.
@@
First, $\langle\mathfrak{C}_{1},H_{b}\rangle=0$ is obvious by definition. Next, by Lemma 2.4, we have $\langle\mathfrak{C}_{1},H_{b}\rangle\equiv\langle\mathfrak{C}^{\prime}_{1},H_{b}\rangle\bmod p$. Let $H^{\prime}$ be a one-component subdiagram of $H_{b}$ with $12n-$ 1 arrows. $H^{\prime}$ must have an arrow connecting two circles of $H_{b}$. Since the tail of the arrow is attached to the circle containing $b$, $H^{\prime}$ is not ascending, and hence $\langle\mathfrak{C}_{2n-1},H_{b}\rangle=0$. ∎
@@
First, a descending diagram is welded equivalent to the trivial one by [16, Proposition 2.2]. Since $\det K$ is invariant under welded equivalence, we conclude that $\det K=1$. ∎
@@
Let $L=L_{1}\cup L_{2}$ be $2$-component mod $2$ almost classical oriented virtual link. If $L$ admits a diagram $D=D_{1}\cup D_{2}$ such that $D_{1}$ is above $D_{2}$ at every crossing between them, then $\det L=0$.
@@
Since $\det B(D_{1})=0$ (see a sentence before [4, Proposition 3.1]), any $(n-1)\times(n-1)$ minor of $B(D)$ is zero. ∎
@@
In the case (I), the contribution of subdiagrams containing $\mathsf{a}$ in $\langle\mathfrak{C}_{2},G_{b}\rangle$ (resp. $\langle\mathfrak{C}_{2},G_{b^{\prime}}\rangle$) is $\varepsilon(r_{+}-r_{-})$ (resp. $\varepsilon(l_{+}-l_{-})$) as drawn in Figure 8. They are the same mod $p$ by Lemma 2.4. Since the contribution of subdiagrams which do not contain $\mathsf{a}$ in $\langle\mathfrak{C}_{2},G_{b}\rangle$ is the same as that of $\langle\mathfrak{C}_{2},G_{b^{\prime}}\rangle$, we conclude that $\langle\mathfrak{C}_{2},G_{b}\rangle\equiv\langle\mathfrak{C}_{2},G_{b^{\prime}}\rangle\bmod 2$.
@@
$\det K$ =±$=\pm\nabla_{\mathrm{asc}}(G)(2)=\pm\sum_{i\geq 0}\langle\mathfrak{C}_{2i},G\rangle 2^{2i}$
±$\equiv\pm 1+\langle\mathfrak{C}_{2},G\rangle 4\equiv\pm 1+4v_{2}(K)\mod 8.$

### v2

@@ 3. Ascending and descending polynomials
We review basic definitions concerning the ascending and descending polynomials according to [7]. An arrow diagram is a based chord diagram with oriented chords (called arrows). A homomorphism $\varphi$ from an arrow diagram $A$ to a based Gauss diagram $G$ is an orientation-preserving homeomorphism of the circle of $A$ to the circle of $G$ which maps the basepoint to the basepoint and induces an injective map of arrows of $A$ to arrows of $G$ respecting the orientation of the arrows.
@@
We recall some terminology from [7, Definitions 4.1 and 4.3]. A chord diagram (or arrow diagram) is said to be $k$-component if the curve obtained from the diagram by taking the parallel double of each chord consists of $k$ circles. Let $A$ be a $1$-component arrow diagram and consider traversing the curve obtained by parallel doubling, starting from the basepoint and following the orientation of the diagram. The diagram $A$ is said to be ascending (resp. descending) if, for each arrow $\mathsf{a}$, the direction in which $\mathsf{a}$ is first traversed is opposite to (resp. the same as) the direction of $\mathsf{a}$.
@@
For a non-negative integer $n$, the Conway combination $\mathfrak{C}_{2n}$ (resp. $\mathfrak{C}^{\prime}_{2n}$) is defined to be the sum of $1$-component ascending (resp. descending) arrow diagrams with $2n$ arrows on a circle. Similarly, $\mathfrak{C}_{2n+1}$ (resp. $\mathfrak{C}^{\prime}_{2n+1}$) is defined to be the sum of $1$-component ascending (resp. descending) diagrams with $2n+1$ arrows on two circles.
@@
The polynomials $\nabla_{\mathrm{asc}}(G)(z)$ and $\nabla_{\mathrm{des}}(G)(z)$ depend on the choice of a basepoint of $G$ in general. For instance, one can observe it for the Gauss diagram in Figure 1. This is because the virtual knot in Figure 1 is not (mod 0) almost classical, which implies that the assumption of Theorems 1.1 and 4.7 is essential in their proofs.
@@
Let $K$ be the (almost classical) virtual knot $6.87548$ appearing in Example 2.12. For the Gauss diagram in [9, Figure 14] (with an arbitrary basepoint), the coefficients of $z^{2}$ in $\nabla_{\mathrm{asc}}(G)(z)$ and in $\nabla_{\mathrm{des}}(G)(z)$ are 2-2. On the other hand, concerning determinant and Arf invariant in [5], one can see $\det(K,F)=5$ and $\operatorname{Arf}(q_{K,F})=1$. Indeed, these are computed from a Seifert matrix
@@
written in [9, Example (6.87548) in Section 6]. See also [5, Remark 3.5].
In [7, Lemma 5.1 and (5.3)], skein relations among pairings are shown for classical links. In their proofs, no property of classical knots is used, and hence we obtain the following lemmas as extensions to virtual links.
@@
Let $_{+}, DD_{-}$, $D_{0}$ be a skein triple drawn in Figure 5 such that $D_{\pm}$ have one component and $D_{0}$ has two components. Let $_{+}, GG_{-}$, $G_{0}$ be the corresponding Gauss diagrams. Then, for $n\geq 1$,
@@
Let $_{+}, DD_{-}$, $D_{0}$ be a skein triple drawn in Figure 5 such that $D_{\pm}$ have two components and $D_{0}$ has one component. Let $_{+}, GG_{-}$, $G_{0}$ be the corresponding Gauss diagrams. Then, for $n\geq 0$,
@@
We extend the notion of warping degree from [19] to based virtual knots. We will use this notion in the proof of Theorem 4.5.
@@
Let $D$ be a diagram of an oriented virtual knot with a basepoint. The warping degree $d(D)$ of $D$ is defined to be the number of crossings at which we first go through the under-arcs when we travel from the basepoint along the orientation of $D$. Moreover, the warping degree $d(G)$ of a based Gauss diagram $G$ is defined to be that of a virtual knot diagram whose Gauss diagram is $G$.
Example 3.9.
Let $D$ be the oriented virtual knot diagram drawn in Figure 4 with the basepoint at the bottom. We then have $d(D)=10$.
A diagram $D$ is said to be descending if $d(D)=0$. It is shown in [17, Proposition 2.2] that a descending diagram is welded equivalent to the trivial one. This fact will be used in the proof of Lemma 4.3.
@@ 4. Main results
First, $\langle\mathfrak{C}_{2n},G_{b}\rangle=0$ is obvious by definition. Next, since $d(G)=0$, every descending subdiagram of $G$ is of the form in Figure 8. This is not $1$-component, and thus $\langle\mathfrak{C}^{\prime}_{2n},G_{b}\rangle=0$. ∎
@@
Let $G_{b}$ be a based Gauss diagram of mod $p$ almost classical oriented virtual knot and let $\mathsf{a}$ be an arrow. Then $\mathsf{a}$ divides the circle into two arcs and suppose that the tail of any arrow intersecting with $\mathsf{a}$ is attached to the arc containing $b$. Let $H_{b}$ be the $2$-component based Gauss diagram obtained by smoothing $G_{b}$ along $\mathsf{a}$. Then, $\langle\mathfrak{C}_{1},H_{b}\rangle\equiv\langle\mathfrak{C}^{\prime}_{1},H_{b}\rangle\equiv 0\bmod p$ and $\langle\mathfrak{C}_{2n-1},H_{b}\rangle=0$ for $n\geq 2$.
@@
First, $\langle\mathfrak{C}_{1},H_{b}\rangle=0$ is obvious by definition. Next, by Lemma 2.5, we have $\langle\mathfrak{C}_{1},H_{b}\rangle\equiv\langle\mathfrak{C}^{\prime}_{1},H_{b}\rangle\bmod p$. Let $H^{\prime}$ be a $1$-component subdiagram of $H_{b}$ with 2n12n-1 arrows. $H^{\prime}$ must have an arrow connecting two circles of $H_{b}$. Since the tail of the arrow is attached to the circle containing $b$, $H^{\prime}$ is not ascending, and hence $\langle\mathfrak{C}_{2n-1},H_{b}\rangle=0$. ∎
@@
First, a descending diagram is welded equivalent to the trivial one by [17, Proposition 2.2]. Since $\det K$ is invariant under welded equivalence, we conclude that $\det K=1$. ∎
@@
Let $L=L_{1}\cup L_{2}$ be a $2$-component mod $2$ almost classical oriented virtual link. If $L$ admits a diagram $D=D_{1}\cup D_{2}$ such that $D_{1}$ is above $D_{2}$ at every real crossing between them, then $\det L=0$.
@@
Since $\det B(D_{1})=0$ (see a sentence before [3, Proposition 3.1]), any $(n-1)\times(n-1)$ minor of $B(D)$ is zero. ∎
@@
In the case (I), the contribution of subdiagrams containing $\mathsf{a}$ in $\langle\mathfrak{C}_{2},G_{b}\rangle$ (resp. $\langle\mathfrak{C}_{2},G_{b^{\prime}}\rangle$) is $\varepsilon(r_{+}-r_{-})$ (resp. $\varepsilon(l_{+}-l_{-})$) as drawn in Figure 9. They are the same modulo $p$ by Lemma 2.5. Since the contribution of subdiagrams which do not contain $\mathsf{a}$ in $\langle\mathfrak{C}_{2},G_{b}\rangle$ is the same as that of $\langle\mathfrak{C}_{2},G_{b^{\prime}}\rangle$, we conclude that $\langle\mathfrak{C}_{2},G_{b}\rangle\equiv\langle\mathfrak{C}_{2},G_{b^{\prime}}\rangle\bmod 2$.
@@
$\det K$ $=\pm\nabla_{\mathrm{asc}}(G)(2)=\pm\sum_{i\geq 0}\langle\mathfrak{C}_{2i},G\rangle 2^{2i}$
$\equiv\pm 1+\langle\mathfrak{C}_{2},G\rangle 4\equiv\pm 1+4v_{2}(K)\mod 8.$